OpenStudy (anonymous):

A room has an area of 456 square feet. One dimension is 5 feet more than the other. Find the dimensions of the room.

OpenStudy (anonymous):

Draw a rectangle. Label the length x and the width (x+5). A = L*W => x(x+5)=456 => x^2 +5x-456=0 => Solve for x using Quad formula. x=19 and x=-24

OpenStudy (anonymous):

So your dimensions are 19 and 24

OpenStudy (anonymous):

thank you!